A teacher realizes that his students' marks on a statistic test are normally distributed with a mean of 62 and a standard deviation of 15. If the teacher wishes to assign minimum marks for grade B is referring to the top 30 per cent of the students' marks in the class, how many marks are required to get a minimum mark for grade B?


1
Expert's answer
2022-04-25T08:20:04-0400

"\mu=62, \sigma=15."

The 70th percentile of a normal distribution is "z=0.524\text{ (find from z-table)};"

"P(Z<0.524)=0.7;\\
P(Z>0.524)=1-0.7=0.3;\\
z=\cfrac{x-\mu}{\sigma};\\"

"x=z\sigma+\mu=0.524\cdot15+62=69.86."

If marks are integers, than the minimum mark for grade B is 70.